y=1/4x+3
1/4x-y=5
(Can you show me the cordnates for the graph?)

Answer by ewatrrr(24785)


Hi
y =1/4x+3
0= 1/4x+3
-3*4 = x (0,3) and (-12,0)
1/4x-y=5 OR y =1/4x - 5 (0,5) and (20,0)